an object of mass 60 kg is raised 12 m and then released. assuming no losses due to air resistance, what is its velocity two-thirds of the way down?


1
Expert's answer
2022-01-30T13:52:00-0500

Answer

Mass of object m=60kg

Raised at height h=12m

So potential energy

U=mgh

"=60\times 9.81\times12\\=7063.2J"

Potential energy at height h'=8m(two-thirds)

"U'=mgh'=60\times9.81\times8"

"\\=4708.8J"

Let kinetic energy is E at this height

Now using energy conservation

"Ui+Ei=Uf+Ef\\7063.2+0=4708.8+E\\E=2354.4J"

So velocity of object

"v=\sqrt{\frac{2E}{m}}\\=\sqrt{\frac{2\times2354.4}{60}}\\=8.86\space m/s"